OpenStudy (anonymous):

Which equation represents a line that is perpendicular to y = -2/3 x + 5 and passes through the point (2,7)

OpenStudy (anonymous):

I got an answer of y= 3/2x+4

OpenStudy (anonymous):

Just want to know if it is correct

OpenStudy (anonymous):

OK. I'll check real quick.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

No, it is not.

OpenStudy (anonymous):

Or is it y= -3/2x+4

OpenStudy (anonymous):

Your first solution intersects the line, but it is not perpendicular. Whatever solution you have, if you are finding a perpendicular line, the slope must be the negative reciprocal of the original line.

OpenStudy (anonymous):

Your second solution is incorrect. I may have a method for you in five minutes or so...

OpenStudy (anonymous):

here is the multiple chooses I have A is y= 3/2x+4 b y=-3/2x+4 c y=3/2x+7 d y= 3/2x+20

OpenStudy (anonymous):

All of the math I have done shows none of the answers to be right; however, I find A is the closest to what I have. Good luck...

OpenStudy (anonymous):

Well thank you then. ^.^ I guess I was right. I will ask my teacher if it is correct. Thank you I appreciate it.